A year hence father will be 5 times as old as his son,two years ago,the father was 3 times as old as his son will be 4 years hence.Find their present ages.
Answer by FrankM(1040)

F+1=5(S+1)
F-2=3(S+4)
5(S+1)-1=3(S+4)+2
5S+4=3S+14
2S=10
S=5 F=29